1. Calculate nominal GDP in 2009 and in 2013 and the percentage increase in nominal GDP from 2009 to 2013. 2. What was the percentage increase in production from 2009 to 2013, and by what percentage did the cost of living rise from 2009 to 2013? 3. In 1988, the average wage rate was $9.45 an hour and in 2008 the average wage rate was $18.00 an hour. The CPI in 1988 was 118.3 and in 2008 it was 215.3. Which real wage rate is higher?


1
Expert's answer
2021-10-19T10:01:56-0400

The base year is 2009, GDP deflator for 2009 will be 1

Real GDP in 2009 was $10 trillion (2009 dollars)

The GDP deflator in 2013 was 112, and real GDP in 2013 was $11 trillion (2009)

a. Real GDP= "\frac{Nominal \; GDP}{GDP \; deflator}"

Thus, nominal GDP in 2009=$10 trillion x 1= $10 trillion

nominal GDP in 2013 = $11 trillion x 1.12= $12.32 trillion

The percentage increase in the nominal GDP between 2009 and 2013 is calculated as

"\frac{Nominal \; GDP \;in \;2013 \;- \;nominal \;GDP\; in \;2009}{Nominal \; GDP \;in \;2009} \times 100"


which is ($12.32 trillion- $10.0 trillion)x100 = 23.2 percent.

b.

The percentage increase in production between 2009 and 2013 is 10 percent.

The percentage increase in the cost of living between 2009 and 2013 is 12 percent.

The percentage change in the cost of living between 2009 and 2013 is calculated as

"\frac{GDP\; price \;index \;in \;2013 \; - \; GDP \;price \;index \;in \;2009}{GDP \;price \;index \;in\; 2009} \times 100"


which is (112-100)/100 x 100 = 12 percent.

c.

Change in the Wages = ($18.00 - $9.45) / $9.45 = 90.5% (90.47)

Change in the CPI or Prices = (215.3 - 118.3) /118.3 = 82% (81.9)

Increase in the real wage = 90.5 - 82 = 8.5%

Real wage today is hgher by 8.5%